OpenStudy (anonymous):

Marcella McCall deposits $969 in a combination of cash and checks. She recalls that the checks she deposited totaled two times the amount of cash she deposited. What amount of checks did she deposit?

OpenStudy (danjs):

gonna set up a system of 2 equations with 2 variables

OpenStudy (danjs):

let C= check H = cash

OpenStudy (danjs):

Total = cash + checks and checks = 2(cash)

OpenStudy (danjs):

make sense so far?

OpenStudy (anonymous):

so i divide 969 by 2?

OpenStudy (danjs):

Deposited 969 dollars in cash and checks 969 = C + H Checks = 2 times the amount of cash C = 2*H

OpenStudy (danjs):

substitute the second into the first 969 = 2*H + H

OpenStudy (anonymous):

whats h ?

OpenStudy (danjs):

H = Cash amount C = Check amount so you have 969 = 3*H H = 323 dollars in cash

OpenStudy (anonymous):

323?

OpenStudy (anonymous):

so why do i divide it by 3?

OpenStudy (danjs):

C= Check H= Cash 969 = C +H and C = 2H Sub second C=2H into first equation 969 = 2H + H 969 = 3H H = 323 323 dollars is cash From second equation, C=2H C=2(323) = 646 646 dollars in Check

OpenStudy (danjs):

As a check, test if the 2 values found agree with both equations.... Total = Checks + Cash 969 = 646 + 323 TRUE, checks out good Checks = 2 times Cash 646 = 2(323) TRUE, Checks out good

OpenStudy (danjs):

you have another one you want to try?

OpenStudy (anonymous):

ooh i understand Thank you (:

OpenStudy (danjs):

no prob, just got to get used to translating sentences into the equations :)